OpenStudy (anonymous):

Solve 4/3 = 12/x 11 9 48/3 8/3

OpenStudy (anonymous):

Cross multiply. \(\dfrac{4}{3} = \dfrac{12}{x}\) 4 * x = ? 3 * 12 = ?

OpenStudy (anonymous):

4x and 36. Wait. You make that into an equation don't you? 4x = 36, right?

OpenStudy (anonymous):

Yep, you got it.

OpenStudy (anonymous):

Now divide 4 to both sides to solve.

OpenStudy (anonymous):

So the answer would be 9.

OpenStudy (anonymous):

Yep, you got it